OpenStudy (anonymous):

what is the lim of x^3-7x/x^3 (x-->0)

OpenStudy (turingtest):

\[\lim_{x\to0}\frac{x^3-7x}{x^3}\]?

OpenStudy (turingtest):

just divide top and bottom by \(x^3\)

OpenStudy (turingtest):

that's all you can really do

OpenStudy (anonymous):

\[\frac{x^3-7 x}{x^3}=1-\frac{7}{x^2} \]As x approaches zero the RHS approaches negative infinity.